All coaches say it takes practice, practice, practice to win. Here's your opportunity to practice to gain the knowledge to win.

Compute the gross earnings for each of the individuals below, rounding each to the nearest cent.

Julie Foudy, an equipment manager, earns $745 for a fluctuating workweek. Overtime work is one-half times the regular rate beyond the 40 hours. Julie worked 7 hours overtime.

  1. Julie's gross earnings = $

Assume Julie's employer computes his pay based on 40-hours per week.

  1. Julie's gross earnings = $

Brandi Chastain, a soccer promotions sales representative, earns an annual base salary of $67,000 and her sales quota is set at $795,000. If she exceeds her quota, she will receive a 7% commission for sales in excess of her quota. For the current year, she has sales of $850,000.

Brandi's gross earnings = $

Hope Solo earns $15.15 per hour and has earned a scoring bonus this week of $45.50. In addition, she worked 6 hours of overtime.

  1. Hope's gross earnings = $
Answer #1

Case 1: Julie Foudy

Total earnings for the week (including overtime) = 745
Total hours worked (including overtime) = 40 + 7 = 47

Let the normal wage rate be = x
Therefore, overtime wage rate is = 1.5x

The information can be represented as the following equation
= (40 * x) + (7 * 1.5x) = 745
= 40x + 10.5x = 745
= 50.5x = 745
= x = $ 14.75 per normal hour
and 14.75 * 1.5 = $ 22.13 per overtime hour

Therefore, Gross earnings = (14.75 * 40) + (22.13 * 7) = $ 745

Case 2: Julie's employer computes his pay based on 40-hours per week

Normal wage rate = $745 / 40 hours = $ 18.625

therefore, overtime wage rate = 18.625 * 1.5 = $ 27.94

Therefore, Gross earnings = (18.625 * 40) + (27.94 * 7) = $ 940.6

Case 3: Brandi Chastain

Current year actual sales = $850,000
Fixed sales quota = $795,000
Therefore, quota exceeded by = $ 55,000
Rate of commission = 7%
Therefore, Commission earned = 3850

Gross annual earnings = annual base salary + commission = 67000 + 3850 = $ 70850

Case 4: Hope Solo

Assumptions : hours per week = 40 hours
Overtime work is one-half times the regular rate beyond the 40 hours

Therefore, Gross earnings = normal wage + overtime wage + weekly bonus

= ($15.5 * 40 hours) + [ ($15.5 * 1.5) * 6 hours ] + $45.5 = $805